# Cubic fourfolds with symplectic automorphisms

### v1

@@ 4.1. The dihedral groups $\mathrm{D}_{10}$ and $\mathrm{D}_{12}$
The action of $\mathrm{D}_{12}$ is given by the follwoing way. A general member of (4.2) is normalized into
@@
by diagonal actions of (×$(\mathbb{C}^{\times})^{6}$. Then $\mathrm{D}_{12}$ is generated by $\mathrm{diag}(-1,-1,1,1,\omega,\omega^{2})$ and a permutation of coordinates by $(12)(56)$.
@@
by diagonal actions of (×$(\mathbb{C}^{\times})^{6}$. Then $\mathrm{D}_{12}$ is generated by $\mathrm{diag}(1,-1,\omega,-\omega,\omega^{2},\omega^{2})$ and a permutation of coordinates by $(13)(24)$.
@@ 4.2. The groups $\mathrm{A}_{3,3}=(\mathrm{S}_{3}\times\mathrm{S}_{3})\cap\mathrm{A}_{6}$
$\begin{array}[]{c|rrrrrr}\rm class&1&\rm 2A&\rm 3A&\rm 3B&\rm 3C&\rm 3D\cr\rm size&1&9&2&2&2&2\cr\hline\cr\chi_{1}&1&1&1&1&1&1\cr\chi_{2}&1&-1&1&1&1&1\cr\chi_{3}&2&0&2&-1&-1&-1\cr\chi_{4}&2&0&-1&2&-1&-1\cr\chi_{5}&2&0&-1&-1&2&-1\cr\chi_{6}&2&0&-1&-1&-1&2\end{array}\quad\begin{array}[]{l}\rm 2A:(12)(45)\\ \rm 3A:(123)\\ \rm 3B:(456)\\ \rm 3C:(123)(456)\\ \rm 3D:(123)(465)\\ \end{array}$
@@
$|\mathrm{Tr}_{\rho}(g)|=\begin{cases}2\quad(\text{the order of $g$ is 2})\\ 0,\,3\quad(\text{the order of $g$ is 3})\end{cases},$
@@
The outer automorphism group $\mathrm{Out}(A_{3,3})$ is isomorphic to $\mathrm{S}_{4}$, and it permutes classes 3A, 3B, $\rm 3A,\ 3B,\ 3C$ and $\rm 3D$. We consider only $\theta_{34}$ and $\theta_{34}\otimes\chi_{2}$ since $\mathrm{Out}(\mathrm{A}_{3,3})$ permutes $\theta_{ij}$ (see Proposition 2.6 (1)).
@@
We see that $\mathrm{Out}(A_{3,3})$ is generated by conjugations by $(12),(13)(24)(36)$, and an automorphism $\sigma$ in Appendix.
A representation $\rho$ affording $\theta_{34}$ is given by permutations of corrdinates $[x_{1}:x_{2}:x_{3}:y_{1}:y_{2}:y_{3}]$, and $V_{3}(\rho)$ is
@@
which is a subspace of (3.6). Although these are stable under permutatoins by $\mathrm{S}_{3,3}$, the actoin of $(12)\in\mathrm{S}_{3,3}$ is non-symplectic. We have
@@
$\chi_{7}+\chi_{10}$ and $\chi_{8}+\chi_{9}$ are not suitable. Let $\rho_{i}$ be a representation affording $\chi_{i}$. We have Molien serieses
@@
(4.10) $\tilde{a}=\begin{bmatrix}a&{\bf 0}\\ {\bf 0}&a\end{bmatrix},\quad\tilde{b}=\begin{bmatrix}b&{\bf 0}\\ {\bf 0}&b\end{bmatrix},\quad\tilde{c}=\begin{bmatrix}c&{\bf 0}\\ {\bf 0}&c\end{bmatrix},\quad$
satisfy symplectic conditions, and $V_{3}(2\rho_{7})$ with projective coordinates $[x_{1}:x_{2}:x_{3}:y_{1}:y_{2}:y_{3}]$ is given by
@@ 4.3. The alternating group $\mathrm{A}_{4}$
Taking projective coordinates $[x_{1}:x_{2}:x_{3}:x_{4}:y_{1}:y_{2}]$, the representation $\rho_{1}$ is given by permutations of $x_{i}$, and it satisfies the symplectic condition. The linear space $V_{3}(\rho_{1})$ is
@@
A general menber of $V_{3}(\rho_{5})$ is smooth since it contains the Fermat cubic. We have
@@
a general menber of $V_{3}(\rho_{5},\chi_{i})$ is not smooth for $i=2,3$.
@@ 5.1. The quaternion group $\mathrm{Q}_{8}$
with usual rules $I^{2}=J^{2}=K^{2}=-1$,  $IJ=-JI=K$. The character table of $\mathrm{Q}_{8}$ is the following:
@@
Since we have $\widehat{\mathrm{Q}_{8}}=\mathrm{C}_{2}\times\mathrm{C}_{2}$, we do not have to consider smi-invariants. If $\rho:\mathrm{Q}_{8}\rightarrow\mathrm{GL}_{6}(\mathbb{C})$ satisfies the symplectic conditions, we have
@@ 5.2. The symmetric group $\mathrm{S}_{4}$
$\begin{array}[]{c|rrrrr}\rm class&1&(12)(34)&(12)&(123)&(1234)\cr\rm size&1&3&6&8&6\cr\hline\cr\chi_{1}&1&1&1&1&1\cr\chi_{2}&1&1&-1&1&-1\cr\chi_{3}&2&2&0&-1&0\cr\chi_{4}&3&-1&-1&0&1\cr\chi_{5}&3&-1&1&0&-1\cr\end{array}$
Since we have $\widehat{\mathrm{S}_{4}}=\mathrm{C}_{2}$, we do not have to consider smi-invariants. If $\rho:\mathrm{S}_{4}\rightarrow\mathrm{GL}_{6}(\mathbb{C})$ satisfies the symplectic condition, we have
@@
and $\rho_{4}$ is not suitable. Taking projective coordinates $[x_{1}:x_{2}:x_{3}:x_{4}:y_{1}:y_{2}]$, the representation $\rho_{1}$ is given by
@@ 6.1. The group $\mathbf{3^{1+4}_{+}:2,\ F_{21}}$ and $\mathbf{QD_{16}}$
An element of order 3 in $G$ is $\mathrm{diag}(1,1,1,\omega,\omega,\omega)$. Moreover, any smooth cubic fourfold with a symplectic automorphism which can be diagonalized as $\mathrm{diag}(1,1,1,\omega,\omega,\omega)$ has this form, and a generic such cubic fourfold has symplectic automorphism group $23^{1+4}_{+}:$ 2.
@@
Extraspecial groups $G=p^{1+2n}_{\pm}$ are non-abelian $p$-groups with center $Z(G)\cong\mathrm{C}_{p}$ and $G/Z(G)\cong\mathrm{C}_{p}^{2n}$. The exponent of $p^{1+2n}_{+}$ is $1$, and that of $p^{1+2n}_{-}$ is $2$. We give an explicit matrix representation of $23^{1+4}:$ 2 in section 7.1.
@@
A general menber in (6.3) is normalized into
@@
by diagonal actions of (×$(\mathbb{C}^{\times})^{6}$. Then $\mathrm{QD}_{16}$ is generated by $a=\mathrm{diag}(1,-1,i,-i,\zeta_{8},\zeta_{8}^{3})$ and a permutation of coordinates by $(34)(56)$.
@@ 6.2. The group $\mathrm{A}_{4,3}=(\mathrm{S}_{4}\times\mathrm{S}_{3})\cap\mathrm{A}_{7}$
Since we have $\widehat{\mathrm{A}_{4,3}}=\mathrm{C}_{2}$, we do not have to consider smi-invariants. A character $\phi$ satisfying conditions
@@
we have $\dim V_{3}(\rho\otimes\chi_{2})=1$ and this is not the case. The representaion $\rho$ is given as follows. Let $[x_{1}:x_{2}:x_{3}:x_{4}:y_{1}:y_{2}:y_{3}]$ be projective coordinates of $\mathbb{P}^{6}$, and $H_{0}$ be a hyperplane
@@ 6.3. The alternating group $\mathbf{\mathrm{A}_{5}}$
Since we have $\widehat{\mathrm{A}_{5}}=1$, we do not have to consider smi-invariants. If $\rho:\mathrm{A}_{5}\rightarrow\mathrm{GL}_{6}(\mathbb{C})$ satisfies the symplectic condition, we have
@@
and $V_{3}(\rho_{3})=0$. For projective coordinates $[x_{1}:\cdots:x_{5}:y]$, the representation $\rho_{1}$ is given as permutations of $x_{i}$, and we have
@@
Although $F\in V_{3}(\rho_{1})$ is stable under permutatoins by $\mathrm{S}_{5}$, the actoin of $(12)\in\mathrm{S}_{5}$ is non-symplectic. Replacing $x_{5}$ and $y$ by $y_{1}$ and $y_{2}$, we see that this is a subspace of (4.20). We have $\dim\mathrm{C}_{\rho_{1}}(\mathrm{A}_{5})=2^{2}+1^{2}=5$ and
@@
The representation $\rho_{2}$ is given by permutations of projective coordinates $[x_{1}:\cdots:x_{6}]$ by “non-standard” $\mathrm{A}_{5}\subset\mathrm{A}_{6}$ generated by $(12345)$ and $(35)(46)$ (see Appendix). Putting $x_{ijk}=x_{i}x_{j}x_{k}$ and
@@
This group $G$ is represented as a (maximal) subgroup of $\mathrm{A}_{6}$ generated by $(123),\ (456),\ (1425)(36)$, and containing $\mathrm{A}_{3,3}=3^{2}:2$ as a subgroup of index 2. We have a Schur cover $G^{*}\cong 3.G$ with GAP-ID (108,15). Firstly, we consider linear representations of $G$ itself. The character table of $G$ is as follows.
@@
for $\chi_{i}$ and $\theta_{34}$ in section 4.2. A representaion $\rho$ affording $\phi_{1}$ is given by permutatioins of coordinates $[x_{1}:x_{2}:x_{3}:y_{1}:y_{2}:y_{3}]$, and $V_{3}(\rho)$ is $(1425)(36)$-invariants of (4.6):
@@
For representations $\rho_{i}$ affording $\eta_{i}$, we have Molien serieses
@@
[…]
@@ 6.5. The groups $\mathbf{\mathrm{S}_{3,3}=\mathrm{S}_{3}\times\mathrm{S}_{3}}$
$\begin{array}[]{c|rrrrrrrrr}\rm class&\rm 1&\rm 2A&\rm 2B&\rm 2C&\rm 3A&\rm 3B&\rm 3C&\rm 6A&\rm 6B\cr\rm size&1&3&3&9&2&2&4&6&6\cr\hline\cr\eta_{11}&1&1&1&1&1&1&1&1&1\cr\eta_{21}&1&-1&1&-1&1&1&1&1&-1\cr\eta_{12}&1&1&-1&-1&1&1&1&-1&1\cr\eta_{22}&1&-1&-1&1&1&1&1&-1&-1\cr\eta_{13}&2&2&0&0&2&-1&-1&0&-1\cr\eta_{23}&2&-2&0&0&2&-1&-1&0&1\cr\eta_{31}&2&0&2&0&-1&2&-1&-1&0\cr\eta_{32}&2&0&-2&0&-1&2&-1&1&0\cr\eta_{33}&4&0&0&0&-2&-2&1&0&0\cr\end{array}\qquad\begin{array}[]{l}\rm 2A=(12)\\ \rm 2B=(45)\\ \rm 2C=(12)(45)\\ \rm 3A=(123)\\ \rm 3B=(456)\\ \rm 3C=(123)(456)\\ \rm 6A=(123)(45)\\ \rm 6B=(12)(456)\\ \end{array}$
@@
Let us consider a representation $\rho_{1}$ affording $\phi_{1}$. Permutations of coordinates $[x_{1}:x_{2}:x_{3}:y_{1}:y_{2}:y_{3}]$ by $\mathrm{A}_{3,3}$ is normalized by an action of a matrix
@@
with eigenvalues $(-1,-1,1,1,1,1)$. These actions generate $\mathrm{S}_{3,3}$ and this gives $\rho_{1}$. By the involution $b_{1}$, the space $V_{3}(\rho)$ in (4.6) is decomposed into the $(+1)$-eigenspace

### v2

@@ 4.1. The dihedral groups $\mathrm{D}_{10}$ and $\mathrm{D}_{12}$
The action of $\mathrm{D}_{12}$ is given by the following way. A general member of (4.2) is normalized into
@@
by diagonal actions of $(\mathbb{C}^{\times})^{6}$. Then $\mathrm{D}_{12}$ is generated by $\mathrm{diag}(-1,-1,1,1,\omega,\omega^{2})$ and a permutation of coordinates by $(12)(56)$.
@@
by diagonal actions of $(\mathbb{C}^{\times})^{6}$. Then $\mathrm{D}_{12}$ is generated by $\mathrm{diag}(1,-1,\omega,-\omega,\omega^{2},\omega^{2})$ and a permutation of coordinates by $(13)(24)$.
@@ 4.2. The groups $\mathrm{A}_{3,3}=(\mathrm{S}_{3}\times\mathrm{S}_{3})\cap\mathrm{A}_{6}$
[…]
@@
[…]
@@
The outer automorphism group $\mathrm{Out}(A_{3,3})$ is isomorphic to $\mathrm{S}_{4}$, and it permutes classes $\rm 3A,\ 3B,\ 3C$ and $\rm 3D$. We consider only $\theta_{34}$ and $\theta_{34}\otimes\chi_{2}$ since $\mathrm{Out}(\mathrm{A}_{3,3})$ permutes $\theta_{ij}$ (see Proposition 2.6 (1)).
@@
We see that $\mathrm{Out}(A_{3,3})$ is generated by conjugations by $(12),(13)(24)(36)$, and an automorphism $\sigma$ in Appendix A.
A representation $\rho$ affording $\theta_{34}$ is given by permutations of coordinates $x_{1}:x_{2}:x_{3}:y_{1}:y_{2}:y_{3}]$, and $V_{3}(\rho)$ is
@@
which is a subspace of (3.6). Although these are stable under permutatoins by $\mathrm{S}_{3,3}$, the action of $(12)\in\mathrm{S}_{3,3}$ is non-symplectic. We have
@@
$\chi_{7}+\chi_{10}$ and $\chi_{8}+\chi_{9}$ are not suitable. Let $\rho_{i}$ be a representation affording $\chi_{i}$. We have Molien series
@@
(4.10) $\tilde{a}=\begin{bmatrix}a&{\bf 0}\\ {\bf 0}&a\end{bmatrix},\quad\tilde{b}=\begin{bmatrix}b&{\bf 0}\\ {\bf 0}&b\end{bmatrix},\quad\tilde{c}=\begin{bmatrix}c&{\bf 0}\\ {\bf 0}&c\end{bmatrix}$$\quad$
satisfy symplectic conditions, and $V_{3}(2\rho_{7})$ with projective coordinates $x_{1}:x_{2}:x_{3}:y_{1}:y_{2}:y_{3}]$ is given by
@@ 4.3. The alternating group $\mathrm{A}_{4}$
Taking projective coordinates $x_{1}:x_{2}:x_{3}:x_{4}:y_{1}:y_{2}]$, the representation $\rho_{1}$ is given by permutations of $x_{i}$, and it satisfies the symplectic condition. The linear space $V_{3}(\rho_{1})$ is
@@
A general member of $V_{3}(\rho_{5})$ is smooth since it contains the Fermat cubic. We have
@@
a general member of $V_{3}(\rho_{5},\chi_{i})$ is not smooth for $i=2,3$.
@@ 5.1. The quaternion group $\mathrm{Q}_{8}$
with usual rules $I^{2}=J^{2}=K^{2}=-1$, IJ=JI=KIJ=-JI=K. The character table of $\mathrm{Q}_{8}$ is the following:
@@
Since we have $\widehat{\mathrm{Q}_{8}}=\mathrm{C}_{2}\times\mathrm{C}_{2}$, we do not have to consider semi-invariants. If $\rho:\mathrm{Q}_{8}\rightarrow\mathrm{GL}_{6}(\mathbb{C})$ satisfies the symplectic conditions, we have
@@ 5.2. The symmetric group $\mathrm{S}_{4}$
$\begin{array}[]{c|rrrrr}\rm class&1&(12)(34)&(12)&(123)&(1234)\cr\rm size&1&3\quad&6&8\quad&6\quad\cr\hline\cr\chi_{1}&1&1\quad&1&1\quad&1\quad\cr\chi_{2}&1&1\quad&-1&1\quad&-1\quad\cr\chi_{3}&2&2\quad&0&-1\quad&0\quad\cr\chi_{4}&3&-1\quad&-1&0\quad&1\quad\cr\chi_{5}&3&-1\quad&1&0\quad&-1\quad\cr\end{array}$
Since we have $\widehat{\mathrm{S}_{4}}=\mathrm{C}_{2}$, we do not have to consider semi-invariants. If $\rho:\mathrm{S}_{4}\rightarrow\mathrm{GL}_{6}(\mathbb{C})$ satisfies the symplectic condition, we have
@@
and $\rho_{4}$ is not suitable. Taking projective coordinates $x_{1}:x_{2}:x_{3}:x_{4}:y_{1}:y_{2}]$, the representation $\rho_{1}$ is given by
@@ 6.1. The group $\mathbf{3^{1+4}_{+}:2,\ F_{21}}$ and $\mathbf{QD_{16}}$
An element of order 3 in $G$ is $\mathrm{diag}(1,1,1,\omega,\omega,\omega)$. Moreover, any smooth cubic fourfold with a symplectic automorphism which can be diagonalized as $\mathrm{diag}(1,1,1,\omega,\omega,\omega)$ has this form, and a generic such cubic fourfold has symplectic automorphism group $3^{1+4}_{+}:2$.
@@
Extraspecial groups $G=p^{1+2n}_{\pm}$ are non-abelian $p$-groups with center $Z(G)\cong\mathrm{C}_{p}$ and $G/Z(G)\cong\mathrm{C}_{p}^{2n}$. The exponent of $p^{1+2n}_{+}$ is $1$, and that of $p^{1+2n}_{-}$ is $2$. We give an explicit matrix representation of $3^{1+4}_{+}2$ in section 7.1.
@@
A general member in (6.3) is normalized into
@@
by diagonal actions of $(\mathbb{C}^{\times})^{6}$. Then $\mathrm{QD}_{16}$ is generated by $a=\mathrm{diag}(1,-1,i,-i,\zeta_{8},\zeta_{8}^{3})$ and a permutation of coordinates by $(34)(56)$.
@@ 6.2. The group $\mathrm{A}_{4,3}=(\mathrm{S}_{4}\times\mathrm{S}_{3})\cap\mathrm{A}_{7}$
Since we have $\widehat{\mathrm{A}_{4,3}}=\mathrm{C}_{2}$, we do not have to consider semi-invariants. A character $\phi$ satisfying conditions
@@
we have $\dim V_{3}(\rho\otimes\chi_{2})=1$ and this is not the case. The representaion $\rho$ is given as follows. Let $x_{1}:x_{2}:x_{3}:x_{4}:y_{1}:y_{2}:y_{3}]$ be projective coordinates of $\mathbb{P}^{6}$, and $H_{0}$ be a hyperplane
@@ 6.3. The alternating group $\mathbf{\mathrm{A}_{5}}$
Since we have $\widehat{\mathrm{A}_{5}}=1$, we do not have to consider semi-invariants. If $\rho:\mathrm{A}_{5}\rightarrow\mathrm{GL}_{6}(\mathbb{C})$ satisfies the symplectic condition, we have
@@
and $V_{3}(\rho_{3})=0$. For projective coordinates $x_{1}:\cdots:x_{5}:y]$, the representation $\rho_{1}$ is given as permutations of $x_{i}$, and we have
@@
Although $F\in V_{3}(\rho_{1})$ is stable under permutatoins by $\mathrm{S}_{5}$, the action of $(12)\in\mathrm{S}_{5}$ is non-symplectic. Replacing $x_{5}$ and $y$ by $y_{1}$ and $y_{2}$, we see that this is a subspace of (4.20). We have $\dim\mathrm{C}_{\rho_{1}}(\mathrm{A}_{5})=2^{2}+1^{2}=5$ and
@@
The representation $\rho_{2}$ is given by permutations of projective coordinates $x_{1}:\cdots:x_{6}]$ by “non-standard” $\mathrm{A}_{5}\subset\mathrm{A}_{6}$ generated by $(12345)$ and $(35)(46)$ (see Appendix A). Putting $x_{ijk}=x_{i}x_{j}x_{k}$ and
@@
This group $G$ is represented as a (maximal) subgroup of $\mathrm{A}_{6}$ generated by (123),(456),(1425)(36)(123),\ (456),\ (1425)(36), and containing $\mathrm{A}_{3,3}=3^{2}:2$ as a subgroup of index 2. We have a Schur cover $G^{*}\cong 3.G$ with GAP-ID (108,15). Firstly, we consider linear representations of $G$ itself. The character table of $G$ is as follows.
@@
for $\chi_{i}$ and $\theta_{34}$ in section 4.2. A representaion $\rho$ affording $\phi_{1}$ is given by permutatioins of coordinates $x_{1}:x_{2}:x_{3}:y_{1}:y_{2}:y_{3}]$, and $V_{3}(\rho)$ is $(1425)(36)$-invariants of (4.6):
@@
For representations $\rho_{i}$ affording $\eta_{i}$, we have Molien series
@@
Eigenvalues of $h$ are 1,1,i1,-1,i, and hence $\tilde{h}$ satisfies the symplectic condition. The group $H\cong 3.A_{3,3}$ generated by $\tilde{a},\tilde{b},\tilde{c}$ in (4.10) is normalized by $\tilde{h}$, and $H$ together with $\tilde{h}$ generate a group $H_{2}\cong 3.(3^{2}.4)$ with GAP-ID (108,15) (we have $\tilde{h}^{2}=\tilde{a}$ and $H_{2}:H]=2$). This gives one of $\rho_{7}+\rho_{8}$ and $\rho_{11}+\rho_{13}$. Moreover we have $\overline{H}=H$ (see (4.9)) and $\overline{h}=h^{3}$, that is, $\overline{H_{2}}=H_{2}$. Therefore we have
@@ 6.5. The groups $\mathbf{\mathrm{S}_{3,3}=\mathrm{S}_{3}\times\mathrm{S}_{3}}$
[…]
@@
Let us consider a representation $\rho_{1}$ affording $\phi_{1}$. Permutations of coordinates $x_{1}:x_{2}:x_{3}:y_{1}:y_{2}:y_{3}]$ by $\mathrm{A}_{3,3}$ is normalized by an action of a matrix
@@
with eigenvalues (1,1,1,1,1,1)(-1,-1,1,1,1,1). These actions generate $\mathrm{S}_{3,3}$ and this gives $\rho_{1}$. By the involution $b_{1}$, the space $V_{3}(\rho)$ in (4.6) is decomposed into the $(+1)$-eigenspace